OpenStudy (anonymous):

x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1 Answer (4, 4, –1) (2, –4, –3) (–3, –1, –5) (–3, 5, –2)

OpenStudy (luffingsails):

Are you using Algebra or Linear Algebra?

OpenStudy (anonymous):

algebra

jhonyy9 (jhonyy9):

x=-16-3y-2z 2x=-15+y-2z so x= (-15+y-2z)/2 -16-3y-2z=(-15+y-2z)/2 -32-6y-4z=-15+y-2z -7y-2z=17 2x-2y-z=1 2x=1+2y+z x=(1+2y+z)/2 (-15+y-2z)/2=(1+2y+z)/2 -15+y-2z=1+2y+z -y-3z=16

jhonyy9 (jhonyy9):

-7y-2z=17 -y-3z=16 so y= -3z-16 than -7(-3z-16)-2z=17 21z+112-2z=17 19z=-95 z= -95/19 so than y= -3(-95/19)-16=285/19 -16 =(285-304)/19 = -19/19 = = -1 y= -1 z= -95/19 - so than x= -3y-2z-16 = -3(-1)-2(-95/19)-16 = 3+190/19-16=3+10-16= = -3 x= -3 y= -1 z= -95/19